OpenStudy (anonymous):

If (x + yi) + 6i = 2, what is x + yi?

OpenStudy (anonymous):

2-6i I think

OpenStudy (anonymous):

See, in case of complex numbers, real part is equal to real part and imaginary part is equal to imaginary part..

OpenStudy (anonymous):

\(x + iy + 6i = 2\)

OpenStudy (anonymous):

The term without i is real and the term with i are imaginary..

OpenStudy (anonymous):

Real on left side is \(x\) and real on right side is 2..

OpenStudy (anonymous):

So, \(x=2\)..

OpenStudy (anonymous):

On left side, imaginary part is : 6i + yi, and on right, there is no imaginary part, or you can say it is 0..

OpenStudy (anonymous):

\(yi + 6i = 0 \\ (y+6)i = 0 \\ y = -6\)

OpenStudy (anonymous):

Therefore, you know x and y values now.. \(x+iy = 2 + i(-6) \\ x + i y = 2 - 6i\)

OpenStudy (anonymous):

Thank you @waterineyes

OpenStudy (anonymous):

\(\dagger\)